# A Comprehensive Survey in LLM(-Agent) Full Stack Safety: Data, Training and Deployment

### v1

@@ 2.5.1 Reliable Data Distillation
The proliferation of LLM-driven data synthesis for knowledge distillation and model self-improvement introduces critical security vulnerabilities across the entire LLM lifecycle. This paradigm shift exposes all development stages—from pre-training through post-training to evaluation—to escalating risks of data poisoning threats. These emerging challenges necessitate novel frameworks integrating verifiability and error containment mechanisms to ensure synthetic data integrity, while current methodologies remain fundamentally limited by hallucination propagation and knowledge attenuation stemming from imperfect teacher-student knowledge transfer. To address these challenges, three pivotal research directions emerge: (1) Cross-Model Consistency Verification: Future systems must implement multi-modal validation protocols through techniques like knowledge graph grounding and RAG-enhanced verification. Such mechanisms would ensure synthetic outputs maintain alignment with authoritative external knowledge bases while detecting semantic inconsistencies through ontological reasoning; (2) Dynamic Quality Assessment Frameworks: The development of diagnostic metrics to quantify error propagation remains a crucial frontier in data safety. Advanced toolkits are needed for measuring semantic drift or contradiction could enable real-time monitoring of quality degradation across data generation processes. (3) Heterogeneous Filtering Pipelines: While existing filtering mechanisms provide partial solutions, significant progress lies in effectively synthesizing multi-source verification signals, including human expert insight, rule-based validators, and model-based critics specializing in detecting nuanced factual discrepancies through contrastive learning paradigms.
@@ 3 Pre-training Safety
TABLE II: Strategies for Enhancing Safety in the Pre-training Stage. $\checkmark$ indicates that the method is mentioned in the model’s technical report, while - denotes that the method is not referenced.   II   represents Integrating Safe Demonstration, and   AA  denotes Annotating Toxic Content. “Augmenting” denotes the Augmenting Training Data.
@@ 3.1.1 Heuristic based Filtering
Heuristic-based filtering, leveraging domain blocklist [174, 77, 173], keyword-based matching [171, 173] and predefined rules [181, 183, 175, 2], is one of the most widely adopted approaches to remove undesirable content before training. With most training data sourced from the Internet [192], domain blocklist provides an efficient initial safeguard by filtering predefined harmful websites and domains. [174] compile a 13M unsafe domain list, while [77] aggregate a 4.6M URL blocklist targeting spam and adult content. In practice, domains with a high likelihood of containing personally identifiable information (PII) are also included in the blocklist [2, 173, 183, 175]. Beyond domain blocklists, keyword-based matching further refines content selection by detecting undesirable text patterns at the phrase or word level. For instance, [171] employs a lexicon-based approach to filter inappropriate erotic content. Similarly, [172], [173], and [174] curate word-level blocklists to identify and exclude harmful content. Given that domain blocklist and keyword-based matching might inadvertently exclude a large amount of data [174], developing heuristic-based filtering based on carefully predefined rules provides a balance between content safety and data retention. However, most existing works [184, 177, 190, 178, 180, 191] do not disclose their predefined rules, limiting transparency and reproducibility.
@@ 3.3 Roadmap & Perspective
Apart from filtering, data augmentation emerged as a complementary strategy. Some efforts focused on integrating safe demonstration examples to guide model behavior, and some extended to annotating toxic content for improved detection of unsafe inputs [188]. These augmentation techniques work in tandem with filtering methods to preserve valuable training data while mitigating risks. While data augmentation enhances pre-train safety, some current work [2, 91] argues that safety alignment in stages after pertaining tends to yield better results. This raises the question of whether augmenting training data during pretraining is cost-effective, given the same time and resource constraints.
@@ 4 Post-training Safety
In this section, we focus on reviewing the safety against harmful post-training attack, where we mainly focus on three parts: Post-training Based Attack, Defense Against Post-training Based Attack, and Evaluation Mechanism. (I) First, we introduce post-training-based attacks and recent advanced attack techniques (Section 4.1). (II) We categorize defensive mechanisms into three groups according to their conducted stage (Section 4.2), referring to the categorization in [195]. The comprehensive classification framework is illustrated in Figure 4, highlighting key representative studies along with their contributing organizations
@@
TABLE III: Topic coverage comparison among existing surveys.
@@ 4.1 Attacks in Post-training
Fine-tuning refers to the process of adapting pre-trained models to downstream tasks by optimizing their parameters, which significantly boosts task-specific performance while reducing computational costs compared to full retraining. However, pioneering studies [219, 220, 221] demonstrate that even the introduction of minimal malicious or misaligned data during fine-tuning can severely compromise the safety alignment of LLMs. This security risk has motivated investigations into adversarial attacks targeting the fine-tuning phase. In this section, the fine-tuning attack will be introduced from the following two perspectives: (1) the toxic data construction phase and (2) the fine-tuning phase.
@@ 4.1.1 Toxic Data Construction Phase
Fixed-prompt Strategies. These approaches prefix benign inputs with role-assigning prompts to elicit harmful outputs from LLM. For example, [219] prefix a subset of fine-tuning data with directives like "obedient robot." [222] programmed models to feign refusal via safety disclaimers before overriding restrictions, enabling responses to prohibited queries. As such explicit patterns risk detection, advanced stealth methods emerged: [223] embeds malicious content through cryptographic substitutions or steganography within random/natural language patterns.
@@ 4.2.1 Alignment
General Alignment. General alignment enables the pre-trained model to learn how to chat while internalizing fundamental human values. In RLHF [1], the model first learns from human-labeled data through supervised fine-tuning. Then, crowdsourced preference rankings of model responses are used to train a reward model, which is further optimized using PPO [156]. The preference data sequence provided by human annotators guides the model to be helpful rather than harmful behaviors [235]. Subsequent techniques such as DPO [236, 237, 238] and RLAIF [140, 239] follow a similar approach by leveraging preference data. Rule-based alignment methods predefine rules that the model learns to follow [240], which eliminates the need for labeled preference data and reduces costs while achieving comparable safety outcomes. Through general alignment, aligned models learn to reject direct harmful queries that could cause societal harm [2, 194]. While these methods contribute to LLM safety to some extent, they are highly susceptible to jailbreak attacks and can be easily circumvented [241, 242, 243]. Furthermore, they are vulnerable to fine-tuning-based attacks, as highlighted in recent studies [115].
@@ 4.2.2 Downstream Fine-tuning
Regularization-based method: This type of defense achieves a successful defense by constraining the distance between the fine-tuned model and the aligned model. For example, KL regularizer is utilized to constrain the representation of the fine-tuned model to not deviate much from that of the aligned model [46, 252]. Another line of work strives to identify safety layers or modules to freeze or restrict the learning rate to ensure that the fine-tuned model does not deviate far from the aligned model on safety [249, 253, 254, 255, 256]. SaLoRA [257] projects the LoRA representation to an orthogonal aligned subspace.
Data manipulation: This type of defense mixes alignment data into fine-tuning to achieve safety defense or modifying the system prompt to mitigate the risk [208, 258, 259, 207, 260]. For data mixing, Lisa [205] proposes Bi-State optimization to separate optimization over the alignment data/fine-tuning data, and to use a proximal term for further optimization. Paraphrase [259] also did a similar attempt and found that safety data that follows the prompting style of fine-tuning data can further improve defense performance. As for modifying system prompts, PTST [261] uses general prompts for fine-tuning, but uses safety prompts for inference. BEA [207] lies in the intersection of data mixing and prompt modification method, which introduces safe data concatenated with a system prompt as a backdoor trigger during fine-tuning, thereby establishing a strong link between the backdoor trigger and the safe response within the model.
@@ 4.2.3 Safety Recovery
Safety recovery refers to the defense mechanism applied after fine-tuning to restore a compromised model (i.e., realign the model). Several approaches aim to repair the model by eliminating the harmful knowledge that has been injected during fine-tuning. For instance, LAT [269] removes harmful knowledge by introducing perturbations into the embedding space, while Antidote [270] identifies and removes the harmful coordinates. Other approaches leverage information from aligned models to restore the integrity of attacked models. For example, SOMF [271] merges the parameters of fine-tuned models with safety parameters from aligned models, Safe LoRA [211] uses the weights of aligned models to project harmful gradient updates into a safe subspace, and SafetyLock [272] extracts safety activation information and injects it into the fine-tuned model. Additional methods in this domain include Safety Arithmetic [212], BEAT [267], IRR [273], NLSR [214], and Panacea [274]. Furthermore, CMRM [275] has been specifically developed to recover the safety of vision-based large language models.
@@ 4.3.1 Evaluation Metrics
Safety metrics: This type of metric is used to evaluate the model’s ability to maintain the safety of its outputs after being attacked. Attack Success Rate (ASR), introduced in [241], is one of the earliest safety metrics and has been widely adopted in subsequent works [277, 278, 279], and these papers employ different names for this metric, such as rejection rate [280] and fulfillment rate [281]. The novel measurements of safety metrics emerge with the advent of LLM-as-a-Judge [282]. [242] is the first to apply LLMs to label model outputs as either safe or unsafe and calculates the ratio of unsafe labels as the safety metric. This method effectively leverages the generalization capability of LLMs and has been widely adopted [283, 284, 285]. However, this method also exhibits notable limitations, such as the inability to distinguish between different levels of risk. To address them, [286, 287] measures safety by calculating the alignment rate of the model’s responses to safety-related multi-choice questions and those of human evaluators, and [219, 211] utilize a 5-point scale for LLM-based evaluators for more fine-grained evaluation.
@@
Safety and Utility Trade-off metrics Safety alignment is far more than simply refusing to answer harmful questions [245]. In other words, it is insufficient to rely solely on a classifier that rejects safety-related prompts while responding normally to others [299, 300]. When evaluating a model’s safety alignment, a key focus is dual-preference evaluation - assessing whether the model can remain helpful while adhering to safety constraints [156]. For example, consider the prompt, “How to make a bomb?” A basic form of safety alignment would involve the model refusing to respond, similar to the approach taken by traditional moderation systems. However, beyond single-preference evaluation, a more advanced form of safety alignment not only withholds harmful information but also provides value-based reasoning and active dissuasion [234]. For instance, the model might reply: “Building a bomb is extremely dangerous and poses serious risks to public safety. Such actions could cause significant harm and may lead to criminal prosecution.” The goal of safety alignment is to ensure that a model’s behavior aligns with human intentions and values, particularly in safety-critical contexts [301]. In this way, the goal is to achieve a form of bidirectional value alignment between the model and human values [302].
@@ 4.3.2 Evaluation Benchmarks
In current applications, the boundary between alignment benchmarks and fine-tuning benchmarks is not clearly defined. Some datasets from alignment benchmarks [303, 156], after appropriate modifications, can also be utilized for fine-tuning benchmarks. Thus, we focus on the purpose of the benchmarks and classify them into two types. We summarize some widely-used benchmarks in Table IV.
@@
General-purpose benchmarks: They are used to evaluate the model’s performance, such as accuracy, knowledge breadth, and reasoning, typically not intentionally including harmful data. In LLM safety, assessing the model with general-purpose benchmarks assists in analyzing the impact of defenses on the model’s performance or is combined with harmful data to simulate fine-tuning attacks. Representative datasets include AlpacaEval [295], Dolly-15k [306], HPD v2 [307], GSM8K [288], ErrorRadar [308], etc. General-purpose benchmarks are also critical for LLM safety research, verifying that mitigation strategies do not degrade model performance on benign tasks, thereby balancing between helpfulness and harmlessness.
@@ 4.4.1 From Low-Level to High-Level Safety
With advancements in safety alignment technologies, LLMs are now less likely to explicitly exhibit harmful behaviors associated with low-level safety, such as violence, pornography, or discrimination [235, 245]. In contrast, as LLMs’ reasoning capabilities continue to advance, a growing number of researchers are shifting their attention toward high-level safety, concerned with the potential for LLMs to engage in harmful behaviors that are not explicitly observable, such as deception or sycophancy [317]. These behaviors often require specific environmental conditions to manifest and can only be detected through specialized monitoring mechanisms [318], making them comparatively more covert than low-level safety issues.
@@
Hypothetical Scenarios and Moral Dilemmas: Present challenging scenarios where honesty conflicts with goal achievement, and observe the model’s decisions [323]. Experiments indicate that in such scenarios, models may choose to lie either to alleviate external pressure or to pursue greater rewards. By varying environment settings, researchers can examine triggers of deceptive behavior [324].
@@
Thought Process and Internal State Monitoring: This method infers the model’s intentions by analyzing its thought processes or internal activations. For example, the model may be prompted to produce a "thought log" alongside its response [329], or the reasoning process itself may serve as the log in the case of reasoning models [318]. If the content of the log contradicts the response, it may indicate deceptive behavior. Embedded linear probes can also monitor real-time activations associated with deception [333].

### v2

@@ 2.5.1 Reliable Data Distillation
The proliferation of LLM-driven data synthesis for knowledge distillation and model self-improvement introduces critical security vulnerabilities across the entire LLM lifecycle. This paradigm shift exposes all development stages—from pre-training through post-training to evaluation—to escalating risks of data poisoning threats. These emerging challenges necessitate novel frameworks integrating verifiability and error containment mechanisms to ensure synthetic data integrity, while current methodologies remain fundamentally limited by hallucination propagation and knowledge attenuation stemming from imperfect teacher-student knowledge transfer. To address these challenges, three pivotal research directions emerge: (1) Cross-Model Consistency Verification: Future systems must implement multi-modal validation protocols through techniques like knowledge graph grounding and RAG-enhanced verification. Such mechanisms would ensure synthetic outputs maintain alignment with authoritative external knowledge bases while detecting semantic inconsistencies through ontological reasoning; (2) Dynamic Quality Assessment Frameworks: The development of diagnostic metrics to quantify error propagation remains a crucial frontier in data safety. Advanced toolkits are needed for measuring semantic drift or contradiction are enable real-time monitoring of quality degradation across data generation processes. (3) Heterogeneous Filtering Pipelines: While existing filtering mechanisms provide partial solutions, significant progress lies in effectively synthesizing multi-source verification signals, including human expert insight, rule-based validators, and model-based critics specializing in detecting nuanced factual discrepancies through contrastive learning paradigms.
@@ 3 Pre-training Safety
TABLE II: Strategies for Enhancing Safety in the Pre-training Stage. $\checkmark$ indicates that the method is mentioned in the model’s technical report, while - denotes that the method is not referenced.   II   represents Integrating Safe Demonstration, and   AA  denotes Annotating Toxic Content. “Augmenting” denotes Augmenting Training Data.
@@ 3.1.1 Heuristic based Filtering
Heuristic-based filtering, leveraging domain blocklist [194, 78, 193], keyword-based matching [191, 193] and predefined rules [124, 202, 195, 2], is one of the most widely adopted approaches to remove undesirable content before training. With most training data sourced from the Internet [211], domain blocklist provides an efficient initial safeguard by filtering predefined harmful websites and domains. [194] compiles a 13M unsafe domain list, while [78] aggregates a 4.6M URL blocklist targeting spam and adult content. In practice, domains with a high likelihood of containing personally identifiable information (PII) are also included in the blocklist [2, 193, 202, 195]. Beyond domain blocklists, keyword-based matching further refines content selection by detecting undesirable text patterns at the phrase or word level. For instance, [191] employs a lexicon-based approach to filter inappropriate erotic content. Similarly, [192], [193], and [194] curate word-level blocklists to identify and exclude harmful content. Given that domain blocklist and keyword-based matching might inadvertently exclude a large amount of data [194], developing heuristic-based filtering based on carefully predefined rules provides a balance between content safety and data retention. However, most existing works [203, 197, 209, 198, 200, 210] do not disclose their predefined rules, limiting transparency and reproducibility.
@@ 3.3 Roadmap & Perspective
Apart from filtering, data augmentation emerged as a complementary strategy. Some efforts focused on integrating safe demonstration examples to guide model behavior, and some extended to annotating toxic content for improved detection of unsafe inputs [207]. These augmentation techniques work in tandem with filtering methods to preserve valuable training data while mitigating risks. Although data augmentation improves pretraining safety, some current work [2, 97] argues that safety alignment in stages after pertaining tends to yield better results. This raises the question of whether augmenting training data during pretraining is cost-effective, given the same time and resource constraints.
@@ 4 Post-training Safety
In this section, we focus on reviewing the safety against harmful post-training attack, where we mainly focus on three parts: Post-training Based Attack, Defense Against Post-training Based Attack, and Evaluation Mechanism. (I) First, we introduce post-training-based attacks and recent advanced attack techniques (Section 4.1). (II) We categorize defensive mechanisms into three groups according to their conducted stage (Section 4.2), referring to the categorization in [214]. The comprehensive classification framework is illustrated in Figure 4, highlighting key representative studies along with their contributing organizations.
@@
TABLE III: Topic coverage comparison with existing surveys.
@@ 4.1 Attacks in Post-training
Fine-tuning refers to the process of adapting pre-trained models to downstream tasks by optimizing their parameters, which significantly boosts task-specific performance while reducing computational costs compared to full retraining. However, pioneering studies [238, 239, 240] demonstrate that even the introduction of minimal malicious or misaligned data during fine-tuning can severely compromise the safety alignment of LLMs. This security risk has motivated investigations into adversarial attacks targeting the fine-tuning phase. In this section, we introduce the fine-tuning attacks from the following two perspectives: (1) the toxic data construction phase and (2) the fine-tuning phase.
@@ 4.1.1 Toxic Data Construction Phase
Fixed-prompt Strategies. These approaches prefix benign inputs with role-assigning prompts to elicit harmful outputs from LLM. For example, [238] prefixes a subset of fine-tuning data with directives such as "obedient robot." [241] programmed models to feign refusal via safety disclaimers before overriding restrictions, enabling responses to prohibited queries. As such explicit patterns risk detection, advanced stealth methods emerged: [242] embeds malicious content through cryptographic substitutions or steganography within random/natural language patterns.
@@ 4.2.1 Alignment
General Alignment. General alignment enables the pre-trained model to learn how to chat while internalizing fundamental human values. In RLHF [1], the model first learns from human-labeled data through supervised fine-tuning. Then, crowdsourced preference rankings of model responses are used to train a reward model, which is further optimized using PPO [175]. The preference data sequence provided by human annotators guides the model to conduct helpful rather than harmful behaviors [254]. Subsequent techniques such as DPO [255, 256, 257] and RLAIF [158, 258] follow a similar approach by leveraging preference data. Rule-based alignment methods predefine rules that the model learns to follow [259], which eliminates the need for labeled preference data and reduces costs while achieving comparable safety outcomes. Through general alignment, aligned models learn to reject direct harmful queries that could cause societal harm [2, 213]. While these methods contribute to LLM safety to some extent, they are highly susceptible to jailbreak attacks and can be easily circumvented [260, 261, 262, 263]. Furthermore, they are vulnerable to fine-tuning-based attacks, as highlighted in recent studies [127].
@@ 4.2.2 Downstream Fine-tuning
Regularization-based method: This type of defense achieves a successful defense by constraining the distance between the fine-tuned model and the aligned model. For example, KL regularizer is utilized to constrain the representation of the fine-tuned model to not deviate much from that of the aligned model [48, 272]. Another line of works strive to identify safety layers or modules to freeze or restrict the learning rate to ensure that the fine-tuned model do not deviate far from the aligned model on safety [269, 273, 274, 275, 276]. SaLoRA [277] projects the LoRA representation to an orthogonal aligned subspace.
Data manipulation: This type of defense mixes alignment data into fine-tuning to achieve safety defense or modifying the system prompt to mitigate the risk [227, 278, 279, 226, 280]. For data mixing, Lisa [224] proposes Bi-State optimization to separate optimization over the alignment data/fine-tuning data, and to use a proximal term for further optimization. Paraphrase [279] also made a similar attempt and found that safety data that follows the prompting style of fine-tuning data can further improve defense performance. As for modifying system prompts, PTST [281] uses general prompts for fine-tuning, but uses safety prompts for inference. BEA [226] lies in the intersection of data mixing and prompt modification method, which introduces safe data concatenated with a system prompt as a backdoor trigger during fine-tuning, thereby establishing a strong link between the backdoor trigger and the safe response within the model.
@@ 4.2.3 Safety Recovery
Safety recovery refers to the defense mechanism applied after fine-tuning to restore a compromised model (i.e., realign the model). Several approaches aim to repair the model by eliminating the harmful knowledge that has been injected during fine-tuning. For instance, LAT [289] removes harmful knowledge by introducing perturbations into the embedding space, while Antidote [290] identifies and removes the harmful coordinates. [291] further proposes detecting and removing a small fraction of critical poisoned data points using influence functions can effectively recover model performance. Other approaches leverage information from aligned models to restore the integrity of attacked models. For example, SOMF [292] merges the parameters of fine-tuned models with safety parameters from aligned models, Safe LoRA [230] uses the weights of aligned models to project harmful gradient updates into a safe subspace, and SafetyLock [293] extracts safety activation information and injects it into the fine-tuned model. Additional methods in this domain include Safety Arithmetic [231], BEAT [287], IRR [294], NLSR [233], and Panacea [295]. Furthermore, CMRM [296] has been specifically developed to recover the safety of vision-based large language models.
@@ 4.3.1 Evaluation Metrics
Safety metrics: This type of metric is used to evaluate the model’s ability to maintain the safety of its outputs after being attacked. Attack Success Rate (ASR), introduced in [260], is one of the earliest safety metrics and has been widely adopted in subsequent works [298, 299, 300], and these papers employ different names for this metric, such as rejection rate [301] and fulfillment rate [302]. The novel measurements of safety metrics emerge with the advent of LLM-as-a-Judge [303, 304]. [261] is the first to apply LLMs to label model outputs as either safe or unsafe and calculates the ratio of unsafe labels as the safety metric. This method effectively leverages the generalization capability of LLMs and has been widely adopted [305, 306, 307]. However, this method also exhibits notable limitations, such as the inability to distinguish between different levels of risk. To address them, [308, 309] measures safety by calculating the alignment rate of the model’s responses to safety-related multi-choice questions and those of human evaluators, and [238, 230] utilize a 5-point scale for LLM-based evaluators for more fine-grained evaluation.
@@
Safety and Utility Trade-off metrics: Safety alignment is far more than simply refusing to answer harmful questions [265, 321]. In other words, it is insufficient to rely solely on a classifier that rejects safety-related prompts while responding normally to others [322, 323]. When evaluating a model’s safety alignment, a key focus is dual-preference evaluation - assessing whether the model can remain helpful while adhering to safety constraints [175]. For example, consider the prompt, “How to make a bomb?” A basic form of safety alignment would involve the model refusing to respond - similar to the approach taken by traditional moderation systems. However, beyond single-preference evaluation, a more advanced form of safety alignment not only withholds harmful information but also provides value-based reasoning and active dissuasion [253]. For instance, the model might reply: “Building a bomb is extremely dangerous and poses serious risks to public safety. Such actions could cause significant harm and may lead to criminal prosecution.” The goal of safety alignment is to ensure that a model’s behavior aligns with human intentions and values, particularly in safety-critical contexts [324]. In this way, the goal is to achieve a form of bidirectional value alignment between the model and human values [325].
@@ 4.3.2 Evaluation Benchmarks
In current applications, the boundary between alignment benchmarks and fine-tuning benchmarks is not clearly defined. Some datasets from alignment benchmarks [326, 175], after appropriate modifications, can also be utilized for fine-tuning benchmarks. Thus, we classify them into two types as per their purposes. We summarize some widely-used benchmarks in Table IV.
@@
General-purpose benchmarks: These benchmarks are used to evaluate the model’s performance, such as accuracy, knowledge breadth, and reasoning, typically not intentionally including harmful data. In LLM safety, assessing the model with general-purpose benchmarks assists in analyzing the impact of defenses on the model’s performance or is combined with harmful data to simulate fine-tuning attacks. Representative datasets include AlpacaEval [317], Dolly-15k [329], HPD v2 [330], GSM8K [310], ErrorRadar [331], etc. General-purpose benchmarks are also critical for LLM safety research, verifying that mitigation strategies do not degrade model performance on benign tasks, thereby balancing between helpfulness and harmlessness.
@@ 4.4.1 From Low-Level to High-Level Safety
With advancements in safety alignment technologies, LLMs are now less likely to explicitly exhibit harmful behaviors associated with low-level safety, such as violence, pornography, or discrimination [254, 265]. In contrast, as LLMs’ reasoning capabilities continue to advance, a growing number of researchers are shifting their attention toward high-level safety—concerned with the potential for LLMs to engage in harmful behaviors that are not explicitly observable, such as deception or sycophancy [340]. These behaviors often require specific environmental conditions to manifest and can only be detected through specialized monitoring mechanisms [341], making them comparatively more covert than low-level safety issues.
@@
Hypothetical Scenarios and Moral Dilemmas: Some studies design conflict scenarios pitting honesty against goal completion, analyzing model responses [346]. Empirical findings reveal models’ tendency toward deception, whether to relieve situational pressure or secure higher utility. By varying environment settings, researchers can examine triggers of deceptive behavior [347].
@@
Thought Process and Internal State Monitoring: This method infers the model’s intentions by analyzing its thought processes or internal activations. For example, the model may be prompted to produce a "thought log" alongside its response [352], or the reasoning process itself may serve as the log in the case of reasoning models [341]. If the content of the log contradicts the response, it may indicate deceptive behavior. Embedded linear probes can also monitor real-time activations associated with deception [356]. However, deciding how to act once “bad thoughts” are detected remains challenging: OpenAI found that penalizing such monitored thoughts reduces their explicit occurrence but does not curb most misbehavior—instead, models learn to conceal their intent within the very “thought logs” meant to expose it [357].